OpenStudy (anonymous):

How can x + 8 = -x + 7 be setup as a system of equations?

OpenStudy (jdoe0001):

\(\large { x + 8 = -x + 7 \implies \begin{cases} {\color{blue}{ y}}=x+8\\ {\color{blue}{ y}}=-x+7 \end{cases}\qquad {\color{blue}{ y}}={\color{blue}{ y}} }\)

OpenStudy (acxbox22):

think about this.... if x+8 is equal to -x+7 if both expressions were equal to another same variable then wouldnt the two equations be the equal itself?

OpenStudy (acxbox22):

i was afk lol technically what @jdoe0001 said